Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Lock acquisition: auxiliary lasers)
mantovani, pinto, masserot, bersanetti, boldrini, tapia, casanueva - 15:55 Friday 19 February 2021 (50836) Print this report
Characterization of the open loop error signal for the HF of ALS
During the shift of thursday 18th, we changed the modulation amplitude at 70kHz as reported in the logbook entry #50829 in order to evaluate if any effects on the noise were present. This modification produced a change in the gain of the global loop, since it affects the amplitude of the error signal. Therefore, to be able to compensate for this change, for each modulation frequency amplitude, the OLTF has been measured at 3888Hz, by injecting a noise line.
We took 7 measurements at different modulation amplitudes, each injection has a duration of 25 sec.
Here are reported the detaild of the measurements, respectively: modulation amplitude; GPS; OLTF_mag @ 3888Hz.

modulation amplitude - GPS- OLTF_mag @ 3888Hz
0.02 1297687040 0.7768
0.025 1297687108 0.9520
0.03 1297687148 1.1334
0.04 1297687197 1.4147
0.05 1297687245 1.6366
0.015 1297687313 0.5877
0.06 1297687372 1.4206

From there we calculated new gains, to match these magnitude points for each measurement (See figure 1.) to be able to obtain a correct computation of the OLTF error signal starting from the closed loop one.
Figure 2 shows the estimated OL error signals, which corresponds to the input noise, together with the closed loop the error signals.
In order to better evaluate the effect of the amplitude of the modulation line the magnitude of the noise has been extracted for different frequency bins and plotted as a function of the modulation amplitude. The results are shown in Figure 3.
It is possible to see that the noise is increasing, by increasing the amplitude of the modulation.
More investigations are needed.
Images attached to this report
Comments to this report:
masserot - 13:01 Monday 22 February 2021 (50847) Print this report

Some more plots with channels at 500KHz or 1MHz fo ALS_NEB for the following cases

  • line 0.02V@70KHz
  • line 0.03V@70KHz
  • line 0.05V@70KHz
  • line 0.06V@70KHz

Control channels:

Monitoring channels

One can see that

  • for the ALS_NEB_PD_GREEN photodiode
    • the noise level increase for the 2 last  70KHs line amplitude  values
    • some strutures appears in the 15KH, 25KHz, 38KHz and 62KHz regions
  • for the monitoring channels, IR or GREEN ones:
    • a peak appears in the 7.5kHz region for  the the 2 last l70KHs line amplitude  values
Images attached to this comment
Search Help
×

Warning

×